Data Table Molarity of Na2S2O3

0.125 M Initial Buret Reading (Na2S2O3)

1.50 mL Final Buret Reading (Na2S2O3) 14.35 mL 


Assume the density of the commercial bleach is 1.08 g/mL. Calculate the average percent by mass of NaClO in the commercial bleach. 


1
Expert's answer
2021-01-11T03:44:53-0500

Solution.

V(sol) = 10.00 mL

"m(sol) = 1.08 \times 10.00 = 10.8 \ g"

"w(NaClO) = \frac{m(NaClO)}{m(sol)} \times 100 \%"

"m(NaClO) = \frac{C(Na2S2O3) \times V(Na2S2O3) \times \frac{M(NaClO)}{2}}{1000} = 0.06 \ g"

w(NaClO) = 0.56 %

Answer:

w(NaClO) = 0.56 %